OpenStudy (anonymous):

solution of x-2/x+4-3

OpenStudy (anonymous):

x-2/x+4 -3=0 multiply evrything by x+4 x-2 -3(x+4)= 0 x-2-3x-12=0 -2x=14 x=-7